y+2=1/2(x-4) whats the standard form?

Answer by ankor@dixie-net.com(22740)

y+2=1/2(x-4) whats the standard form?
get rid of the fraction multiply both sides by 2
2(y+2) = x - 4
2y + 4 = x - 4
-x + 2y = -4 - 4
-x + 2y = -8 is the standard form